OpenStudy (clara1223):

Find the equation of the osculating circle to y=(1/2)x^2 at the point (1,1/2).

OpenStudy (clara1223):

In class we found that y'=x and y''=1, curvature=\[\frac{ y'' }{ (1+y'^{2})^{\frac{ 3 }{ 2 }} }\] so curvature=\[\frac{ 1 }{ 2^{\frac{ 3 }{ 2 }} }\] to find the center of the circle we did \[(1,\frac{ 1 }{ 2 })+2^{\frac{ 3 }{ 2 }}(-\frac{ 1 }{ \sqrt{2}}+\frac{ 1 }{ \sqrt{2} })\] simplifying: \[(-1,\frac{ 5 }{ 2 })\] so the equation of the circle is \[(x+1)^{2}+(y-\frac{ 5 }{ 2 })^{2}=8\] What I am confused about is where the \[(-\frac{ 1 }{ \sqrt{2} }+\frac{ 1 }{ \sqrt{2} })\] came from.

OpenStudy (dallascowboys88):

a

OpenStudy (vanesaretana):

The correct answer is the 1st one.

OpenStudy (dumbcow):

@clara1223 the \[(-\frac{1}{\sqrt{2}}, \frac{1}{\sqrt{2}})\] comes from using polar form to obtain "x" distance and "y" distance from point to center of circle Note from polar form that \[x = r \cos \theta\] \[y = r \sin \theta\] where "r" is radius Also \[-\frac{1}{y'} = \tan \theta\] Where tan is slope of normal line and y' is slope of tangent line Using your example: y' = x , at point (1,1/2) ----> y' = 1 \[\tan \theta = -1 \rightarrow \theta = \frac{3\pi}{4}\] \[\cos \theta = -\frac{1}{\sqrt{2}}, \sin \theta = \frac{1}{\sqrt{2}}\] \[r = 2^{3/2} \] Therefore the center of circle is: \[(x,y) + r(\cos \theta , \sin \theta)\] \[(1, \frac{1}{2}) + 2^{3/2}(-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}})\] Hope that explains it for you.

likeabossssssss (likeabossssssss):

@dallascowboys88 and @Vanesaretana no dirent answer plz

likeabossssssss (likeabossssssss):

that is cheating